40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1316     Entry time: Tue Feb 17 05:20:11 2009
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: Locking 
Since we excluded *.snap and *.req files from the svn control in the medm directory and these were not restored by the svn co, the burt part of the align/mis-align scripts were not working correctly this evening. So I recreated .req files and cooked up some mis-aligned .snap files.
After some cut-and-try work, I was able to run the dither alignment scripts fine.

Due to the above mentioned delay, the locking work started around midnight.

Tonight, the DD hand-off was not robust. I spent sometime to optimize this.
After the optimization, the locking proceeded to the DC CARM/DARM control state stably.
The CARM->MCL hand-off failed because the LSC-MC offset button was off.
I added a line to turn on the button in the ontoMCL script.
Today, the offloadMCF script worked fine.

Next, the cm_step script stumbled on the "ENGAGERIZING" of the AO path.
I got a hunch that the AO path might not be connected to the MC board.
Indeed, OMC_OSC_FM was connected to the IN2 of the MC board. Looks like it was used for the optimization of the modulation frequencies.
Probably I had the hunch because I did it Smile

I was able to increase the arm power up to 3.9.
The script failed when it tried to switch the CARM signal from TR_DC to SPOB_DC.
I haven't tackled on this issue yet.
ELOG V3.1.3-